OpenStudy (anonymous):

A Norman window has the shape of a rectangle capped by a semicircle. What is the length of the base of a Norman window of maximum area if the perimeter of the window equals 7?

OpenStudy (anonymous):

Let r = radius of semi-circle (half the base). Let h = height of rectangle Perimeter = 9 = 2r + 2h + pi r [base + sides + arc] r (pi + 2) + 2h = 9 2h = 9 - r (pi + 2) h = 9/2 - r/2 (pi + 2) Area = 2r * h + pi r^2 / 2 = 2r (9/2 - r/2 (pi + 2)) + pi r^2 / 2= 9 r - r^2 (pi + 2) + pi r^2 / 2 = 9 r - r^2 (pi + 2 - p/2) = 9 r - r^2 (pi/2 + 2) Derivative is: 9 - r (pi + 4) Setting that to 0 to find maximum: r = 9 / (pi + 4) = 9 / (4 + pi) = 1.2602231 For this value, call it 1.26, arc = pi * 1.26 = 3.9584 height = ( 9 - 2.52 - 3.96 ) / 2 = 1.26 So the rectangular part consists of 2 squares 1.26 x 1.26 The base is 2.52 (or 18 / (4 + pi) to be exact). Maximum area problems of this kind *always* turn out to have the dimensions "most evenly divided" into parts of the perimeter (square, circle, other similar shape).

OpenStudy (anonymous):

I was looking at those problems and some of them I don't understand where they get of the number so I was completely lost.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you know what a norman window is?

OpenStudy (anonymous):